OpenStudy (anonymous):

Find f(2) if f(x) = (x + 1)2. 5 6 9

OpenStudy (anonymous):

\[f(x)=(x+1)^2, f(2)=(2+1)^2\]

OpenStudy (nurali):

f(x) = (x + 1)2 f(2)=(2+1)^2 f(2)=3^2 f(2)=9